OpenStudy (anonymous):

Find f(2)(0) if f(x)= x^2-1/x^2+2x+1. Is f(0) a relative maximum or a relative minimum of f?

OpenStudy (freckles):

We need to first find the first derivative. I assume you mean (x^2-1)/(x^2+2x+1) Though I could be wrong and you actually mean x^2-1/x^2+2x+1

OpenStudy (anonymous):

youre correct

OpenStudy (freckles):

First of all we could make this prettier by canceling a factor from numerator in denominator of (x^2-1)/(x^2+2x+1) we don't have to do this but I think it will make our future with this problem much better.

OpenStudy (anonymous):

we cancel x+1?

OpenStudy (freckles):

yep

OpenStudy (freckles):

\[f(x)=\frac{x-1}{x+1}\] Doesn't this look much easier to find the derivative of? So what is f' You need to use what rule?

OpenStudy (anonymous):

quotient deriv

OpenStudy (freckles):

yep

OpenStudy (anonymous):

so its (x+1)(x) - (x-1)(x) / (x+1)^2

OpenStudy (freckles):

\[f'(x)=\frac{(x-1)'(x+1)-(x+1)'(x-1)}{(x+1)^2} \] I think you mean derivative of either (x-1) or (x+1) is 1 not x.

OpenStudy (anonymous):

ops!

OpenStudy (freckles):

So simplify or f' then we will continue to go ahead and find the second derivative

OpenStudy (anonymous):

im not sure how you simplify this...

OpenStudy (freckles):

You don't know how to simplify (x+1)-(x-1) ?

OpenStudy (anonymous):

oh but wat about the denominator?

OpenStudy (freckles):

It is already simplified

OpenStudy (anonymous):

i got 2

OpenStudy (freckles):

\[f'(x)=\frac{2}{(x+1)^2} \]

OpenStudy (freckles):

now find the second derivative

OpenStudy (anonymous):

(x+1)^2 (0) - (2) 2(x+1) / (x+1)^4

OpenStudy (anonymous):

i think

OpenStudy (freckles):

\[f''(x)=\frac{(2)'(x+1)^2-2((x+1)^2)'}{(x+1)^4}\] You did great. Now simplify the top.

OpenStudy (anonymous):

4x+4

OpenStudy (anonymous):

or is it -4x-4

OpenStudy (freckles):

on top you have -2(2)(x+1)

OpenStudy (freckles):

-4(x+1)/(x+1)^4 you can actually reduce this fraction

OpenStudy (anonymous):

-4/(x+1)^3

OpenStudy (freckles):

yep that is the second derivative

OpenStudy (freckles):

now surely they don't mean "is f(0) a rel min or max?" because x=0 is not even a critical number

OpenStudy (anonymous):